OpenStudy (anonymous):

Rationalize the denominator and simplify. All variables represent positive real numbers. square root 11 / square root 33 - square root 5

OpenStudy (radar):

\[\sqrt{11}\div(\sqrt{33}-\sqrt{5})\]

OpenStudy (radar):

Is that the problem the denominator consist of two terms

OpenStudy (radar):

assuming that is the case multiply the numerator and the denominator by the conjugate of the denominator.

OpenStudy (radar):

\[\sqrt{11}(\sqrt{33}+\sqrt{5})\div(\sqrt{33}-\sqrt{5})(\sqrt{33}+\sqrt{5})\]

OpenStudy (radar):

\[(\sqrt{11}\sqrt{33}+\sqrt{11}\sqrt{5})\div(33-5)\]

OpenStudy (anonymous):

ok what do you do next

OpenStudy (radar):

\[(11\times \sqrt{3}+\sqrt{11}\sqrt{5})\div28\]

OpenStudy (radar):

Think that is far as I can go with it. Doesn't seem to simplify any further! But it sure don't look simple. LOL

OpenStudy (radar):

\[(11\times \sqrt{3}+\sqrt{55})\div28\]

OpenStudy (radar):

A little bit simpler